OpenStudy (anonymous):

solve without using calculator 5ln(x-2)-2=28 help please, thank you

OpenStudy (anonymous):

5ln(x-2) = 30

OpenStudy (anonymous):

ln(x-2) = 6

OpenStudy (anonymous):

ln or log ?

OpenStudy (mimi_x3):

\( 5ln(x-2) = 30 \) \(ln(x-2) = 6\) \(log_e(x-2) = 6\) \((x-2) = e^{6}\) Are you able to do it now?

OpenStudy (anonymous):

ln

OpenStudy (anonymous):

im stuck at the (x-2)=e^6

OpenStudy (mimi_x3):

bring the \(2\) to the RHS

OpenStudy (anonymous):

i know the answer is e^6 +2 but i dont know where the =2 comes from

OpenStudy (anonymous):

+2

OpenStudy (mimi_x3):

Because when you bring the \(-2\) to the RHS it's a positive

OpenStudy (mimi_x3):

I don't know the correct way to say that lol

OpenStudy (anonymous):

alright can you walk me through the part after ln(x-2)=6

OpenStudy (mimi_x3):

Well you know that \(ln\) can also be written as \(log_e\) right?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

so you get e^6= (x-2) right

OpenStudy (mimi_x3):

yeah \((x-2) = e^{6}\)

OpenStudy (anonymous):

so what do you do next

OpenStudy (mimi_x3):

well i just told you before :)

OpenStudy (anonymous):

bring the 2 to the RHS?

OpenStudy (anonymous):

ooooohhhhhhhh xD duh x = e^6 + 2

OpenStudy (mimi_x3):

yes; you dont really bring it was called something else which forgot. just simply drag it to RHS. But there is a better name for it lol

OpenStudy (anonymous):

i forgot i was solving for x , silly me

OpenStudy (anonymous):

i more last question what do you do for e^x-3 = 16

OpenStudy (mimi_x3):

\[ e^x - 3 =16\] What have you tried?

OpenStudy (anonymous):

e^x-3 + 4 =20

OpenStudy (mimi_x3):

Um, that is not right..

OpenStudy (mimi_x3):

drag the \(-3) to the rhs

OpenStudy (anonymous):

but its an exponent x-3 is the exp. for e

OpenStudy (mimi_x3):

\[ e^x = 16+ 3\]

OpenStudy (mimi_x3):

\[ e^x = 19\]

OpenStudy (mimi_x3):

Take ln both sides \[xlne = ln19\]

OpenStudy (anonymous):

the answer is ln(16)+3

OpenStudy (mimi_x3):

Wolfram Alpha agrees with me http://www.wolframalpha.com/input/?i=+e%5Ex-3+%3D+16

OpenStudy (anonymous):

my teacger myst be wring though . she gave me that answer

OpenStudy (anonymous):

my teacher was wron. she gave me that answer

OpenStudy (mimi_x3):

well there is a low chance that wolfram alpha is wrong

OpenStudy (anonymous):

oh, well, it looks like we have a paranthesis confusion here. when writing e^x-3 = 16, it's important that you write the parantheses aroudnthe (x-3), else we can't know that the whole thing is the exponant. So we have this right? \[e ^{x-3} = 16\]